OpenStudy (anonymous):

integral of sqrt 1+81x

OpenStudy (anonymous):

mimi...no lmao just no

OpenStudy (anonymous):

its sqrt (1+81x)

OpenStudy (anonymous):

You can do a u sub where u=1+81x and du=81dx: \[\frac{1}{81}\int\limits \sqrt{u} \; du\]

OpenStudy (mimi_x3):

lols, sorry, i didn't know that its sqrt(1+81x) put it in brackets next time, i misunderstood it.

OpenStudy (mimi_x3):

 = 2/243(1+81x)^3/2